# Flying Bird Object Detection Algorithm in Surveillance Video

### v1

@@ Abstract
Aiming at the characteristics of the flying bird object in surveillance video, such as the single frame image feature is not obvious, the size is small in most cases, and asymmetric, this paper proposes a Flying Bird Object Detection method for Surveillance Video (FBOD-SV). Firstly, a new feature aggregation module, the Correlation Attention Feature Aggregation (Co-Attention-FA) module, is designed to aggregate the features of the flying bird object according to the bird object’s correlation on multiple consecutive frames of images. Secondly, a Flying Bird Object Detection Network (FBOD-Net) with down-sampling and then up-sampling is designed, which uses a large feature layer that fuses fine spatial information and large receptive field information to detect special multi-scale (mostly small-scale) bird objects. Finally, the SimOTA dynamic label allocation method is applied to One-Category object detection, and the SimOTA-OC dynamic label strategy is proposed to solve the difficult problem of label allocation caused by irregular flying bird objects. In this paper, the algorithm’s performance is verified by the experimental data set of the surveillance video of the flying bird object of the traction substation. The experimental results show that the surveillance video flying bird object detection method proposed in this paper effectively improves the detection performance of flying bird objects. This project is publicly available https://github.com/Ziwei89/FBOD.
@@ I Introduction
BIRD detection plays an extremely important role in ensuring safety and alleviating conflicts between humans and animals and has attracted more and more researchers’ attention in recent years. At present, the method of detecting bird objects by radar is widely used [1, 2], but the radar equipment has the disadvantages of being large in volume, expensive, and having poor visual effects. With the development of computer vision, deep learning, and other technologies, there are more and more studies on using cameras to detect various objects [3, 4, 5, 6]. We are working on the real-time detection of flying birds using surveillance cameras. However, there are three main challenges for flying birds in surveillance video.
@@
    The characteristics of the single frame image of the flying bird object in the surveillance video are not obvious. In most cases, the flying bird object blends into the background, making it challenging to distinguish from the environment, even after careful examination, as shown in Fig. 1LABEL:sub@obvious_bird_fig left.
@@
    Most flying bird objects in the surveillance video are small. As Fig. 1LABEL:sub@special_scale_bird_fig, demonstrates, when birds are far from the camera, they occupy fewer pixels in the video frame, categorizing them as small objects. Conversely, when birds are closer to the camera, they occupy more pixels and can be considered large objects. However, birds are often far away from the camera, resulting in them being predominantly classified as small objects.
@@
    In most cases, the flying birds in the surveillance video are not symmetrical and irregular in the bounding box. For example, in some cases, the body and tail of the flying bird in the bounding box are half and half, and half of the tail has more background pixels. When the flying bird flings its wings, the part without wings contains more background pixels, as shown in Fig. 1LABEL:sub@irregularity_scale_bird_fig.
@@
Figure 1: Characteristics of flying bird objects in surveillance videos. (a) On the right, there is a small bird with weak features. Left is a screenshot of the bird on five consecutive frames. (b) Simplified diagram of the distribution of flying birds in surveillance video. Birds are evenly distributed in the surveillance area (A, B, and C areas: birds in area A are large targets, birds in area B are general targets, and birds in area C are small targets). However, the space of area C is much larger than that of areas A and B, so birds are mostly small objects. (c) In the right image, the green box is the bounding box, and the red point is the middle of the bounding box. Due to the shape of the bird object itself, it is not regular in the bounding box in most cases.
@@
Some papers have specifically studied the detection methods of flying bird objects. For example, literature [23] proposed a skeleton-based bird object detection method, which describes the motion information of birds through a set of key postures. Skeleton features are selected as relatively fixed and common features in bird posture appearance to overcome the variability of bird posture. Another example is Glances and Stare Detection(GSD) [24], which simulates the characteristics of human observation of small moving objects. Firstly, the whole image is glanced at, and the places containing objects are carefully observed. Specifically, the region proposal sequence obtained by localization is first used to construct the candidate region sequence, and then the classification method is used to classify the candidate region sequence, and the classification results are fed back to localization to guide the generation of new region proposals. By alternating localization and classification, the region proposal shrinks until the bird object is locked. Although these methods consider some characteristics of the flying birds, they still require the flying birds to have a relatively clear appearance, and their methods are still difficult to apply to the actual surveillance video.
@@
Aiming at the problem that the features of the single frame image of the flying bird object in the surveillance video are not obvious, our previous work [25] used the Convolutional Long Short-Term Memory network (ConvLSTM) to aggregate the spatio-temporal information of adjacent multiple frames of flying bird objects at the input of the model to enhance the characteristics of flying bird objects, to improve the detection rate of flying bird objects. However, ConvLSTM runs slowly, and the information aggregation effect is not good.
Although the characteristics of the single frame image of the flying bird object in the surveillance video are not obvious in some cases, the object can still be found by observing the continuous multiple frames of images, as shown in Fig. 1LABEL:sub@obvious_bird_fig right. Therefore, in some cases, the information from other image frames can be used to observe the object on one frame because the bird object has a correlated relationship between frames. In this paper, we take advantage of this correlation relationship and propose a new method to aggregate the information of consecutive multiple frames of images, the Correlation Attention Feature Aggregation (Co-Attention-FA) method, which uses the correlation relationship of flying bird objects on consecutive multiple frames of images to ignore redundant background information and focus on the region containing the flying bird object.
Aiming at the special multi-scale property of flying birds in surveillance videos (most of them belong to small objects), a Flying Bird Object Detection Network (FBOD-Net) is designed, which first down-sampling and then up-sampling. A large feature layer that combines fine spatial information and large receptive field information is used to detect flying birds in surveillance videos. The large feature layer has certain advantages for detecting small-scale size bird objects because it has fine spatial information (shallow, high-resolution feature layer) and highly abstract semantic information (deep feature layer). Due to the fusion of large receptive field information, the large feature layer also has the ability to detect large-scale size flying bird objects  [^1] FBOD-Net is trained on datasets dominated by small objects with occasional large objects and can detect large objects. However, the training process may not converge if datasets contain mainly large objects..
Aiming at the irregular characteristics of the bird object in the bounding box in the surveillance video, the SimOTA (SimOTA-OC) dynamic label strategy for One-Category object detection was proposed. When the SimOTA [16] dynamic label allocation method is applied to the bird object detection (one-category object detection), only the IOU between the Ground Truth (GT) box and the predicted box is used to realize the dynamic allocation of labels, which solves the problem of inaccurate label allocation caused by the irregular bird object.
@@
The remainder of this paper is structured as follows: Section II presents work related to this paper. Section III describes in detail the method of flying bird object detection in surveillance video. In Section IV, the ablation and comparison experiments of the proposed algorithm are carried out. Section V concludes our work.
@@ II-A Related work on Feature aggregation
The related research work of feature aggregation mainly deals with video object detection. When the features of the object to be detected on some video frames are not obvious (appearance change, occlusion, motion blur), the feature information of other adjacent frames can be aggregated to enhance the features of the object on these frames in the feature extraction stage.
Zhu Xizhou et al. [26, 27, 28, 29] and Wang Shiyao et al. [30] used optical flow to propagate features and then aggregated the features of the current frame and the propagated features to enhance the expression ability of the current frame to improve the detection accuracy of video objects. Yongyi Lu et al. [31] and Mason Liu et al. [32] used the characteristics of LSTM to realize the aggregation or association of feature information of different frames to realize the detection of video objects. Hao Luo et al. [33] aggregate features that enhance each candidate by modeling the semantic and spatio-temporal relationships between object candidates within the same frame and between adjacent frames. Haiping Wu et al. [34] proposed a Sequent-level Semantic Aggregation (SELSA) method, which randomly samples some image frames from the whole video and aggregates the object candidates of the current frame with the weight of the distance between the object candidates to enhance the characteristics of the object candidates of the current frame. Tao Gong et al. [35] aggregated the ROI features of the current frame and the most similar ROI features of other frames to obtain the Temporal ROI features so that the Temporal ROI features contained the temporal domain information of the object to be detected in the whole video. Liang Han et al. [36] combined class-aware pixel-level feature fusion and class-aware instance-level feature fusion to improve the detection accuracy of video objects. Chao Xu et al. [37] introduced a multi-level spatio-temporal feature aggregation framework to fully use frame-level, proposal-level, and ince-level information under a unified framework to improve the detection accuracy of video objects. In addition, Hanming Deng et al. [38] and Masato Fujitake et al. [39] designed a unique external memory to store the features of other frames and then fused the features of the current frame to complete feature aggregation.
In the previous work [25], we pointed out that the above method of feature aggregation, whose aggregation operation is carried out in the intermediate feature layer, is unsuitable for the flying bird object in the surveillance video. Because the appearance features of most of the bird objects in the surveillance video are not particularly rich in any single frame, and the object size is small, the features of the object are easy to lose when extracting the features of a single frame image. After feature extraction, aggregation is easy to introduce some wrong information. Therefore, in the previous work [25], we used the ConvLSTM to aggregate the spatio-temporal information of adjacent multiple frames of flying bird objects at the input of the model to enhance the characteristics of flying bird objects, to improve the detection rate of the flying bird object. However, ConvLSTM runs slowly, and the information aggregation effect is not good.
@@ III The Proposed FBOD-SV
Fig. 2 shows the summary diagram of the proposed Flying Bird Object Detection method for Surveillance Video (FBOD-SV), which mainly includes four parts, namely, the Co-Attention-FA unit, the FBOD-Net model, the SimOTA-OC label assignment unit, and the model training unit. Among them, the Co-Attention-FA unit aggregates the features of the flying bird object on consecutive frames of images. The FBOD-Net model uses the aggregated features of the flying bird object to predict the location information of the flying bird object. The SimOTA-OC label assignment unit dynamically allocated positive samples according to the model prediction results during model training. The multi-task loss function is used in the model training unit to train the flying bird object detection model. Next, the Co-Attention-FA unit (III-A), the FBOD-Net (III-B), the SimOTA-OC dynamic label assignment method (III-C), and the loss function (III-D) will be introduced, respectively.
@@ III-A The Co-Attention-FA
The diagram of the Co-Attention-FA module is shown in Fig. 3. Firstly, the consecutive frames of images $\left\{\text{x}_{1},\text{x}_{2},...,\text{x}_{n}\right\}$ are concated into a multi-channel feature matrix X, prepared for the information correlation between frames. Then, the information of consecutive frames is correlated and fused by convolution to obtain the feature matrix $\text{F}_{1}$. At this time, $\text{F}_{1}$ already has the correlation information between frames.
@@
Next is a residual structure. Inside the residual structure, $\text{F}_{1}$ is convolved to get $\text{F}_{2}$, and the correlation information between frames is used to make $\text{F}_{2}$ focus on the bird object. Then $\text{F}_{2}$ is operated by the Sigmoid function to obtain $\text{F}_{3}$ to make the focusing information explicit, $\text{F}_{4}$ is obtained by dot multiplication of $\text{F}_{3}$ and $\text{F}_{2}$ to realize spatial attention.
On the outside of the residual structure, $\text{F}_{4}$ is superimposed on $\text{F}_{1}$. The purpose of such processing is to highlight the characteristics of the flying bird object while retaining the necessary background information. Finally, the aggregated feature $\text{F}_{6}$ is output by a convolution operation.
Fig. 4 shows the schematic of aggregating continuous video image information using the Co-Attention-FA module.
Figure 4: To aggregate continuous video image information using the Co-Attention-FA module.
@@ III-B The FBOD-Net
Deep Neural Network (DNN) is generally useful for detecting large objects by down-sampling multiple times to obtain a large receptive field and strong semantic information. However, multiple down-samplings result in low resolution of the feature map, and small objects are easily lost. The shallow feature map of the neural network has high resolution and rich spatial detail information, which can be used to detect small objects. Therefore, for objects with multi-scale attributes, the deep small feature maps and the shallow large feature maps are generally combined to detect multi-scale objects.
Since the flying birds in surveillance videos have special multi-scale properties (i.e., most are small objects), we adopt a large feature map to predict the flying birds in surveillance videos. However, the shallow large feature maps do not have a large receptive field, which is not conducive to detecting large objects. At the same time, the semantic information of the shallow large feature maps is weak, which is not conducive to object recognition. Therefore, we fuse the information of the deep small feature map of DNN into the large feature map so that the large feature map has both delicate spatial information and abstract high-level semantic information.
Specifically, we adopt a network structure with first down-sampling and then up-sampling to fully fuse the shallow and deep feature map information, as shown in Fig. 5. Firstly, CSPDarkNet53 [49] was used to extract the down-sampling style features of the feature aggregated image, and a series of feature maps ($\text{C}_{1}$, $\text{C}_{2}$, $\text{C}_{3}$,$\text{C}_{4}$, $\text{C}_{5}$ in Fig. 5) were obtained. Then, the convolution + Pyramid Pooling (SPP)+ convolution operation is performed on $\text{C}_{5}$ to obtain the feature map $\text{P}_{5}$. $\text{P}_{5}$ is upsampled (the bilinear interpolation method is used in this paper) and then fused with $\text{C}_{4}$ (a schematic of the fusion operation is shown in Fig. 6) to obtain the feature map $\text{P}_{4}$. The feature layers $\text{P}_{4}$, $\text{P}_{3}$, and $\text{P}_{2}$ perform the same operation as $\text{P}_{5}$ (up-sampling and then fusion with the shallow feature map of the same scale) and finally obtain the feature map $\text{P}_{1}$. Finally, the feature map $\text{P}_{1}$ undergoes two different convolution operations simultaneously to obtain the feature map of prediction confidence and the feature map of regression flying bird object position.

### v3

# A Flying Bird Object Detection Method for Surveillance Video

@@ Abstract
Aiming at the specific characteristics of flying bird objects in surveillance video, such as the typically non-obvious features in single-frame images, small size in most instances, and asymmetric shapes, this paper proposes a Flying Bird Object Detection method for Surveillance Video (FBOD-SV). Firstly, a new feature aggregation module, the Correlation Attention Feature Aggregation (Co-Attention-FA) module, is designed to aggregate the features of the flying bird object according to the bird object’s correlation on multiple consecutive frames of images. Secondly, a Flying Bird Object Detection Network (FBOD-Net) with down-sampling followed by up-sampling is designed, which utilizes a large feature layer that fuses fine spatial information and large receptive field information to detect special multi-scale (mostly small-scale) bird objects. Finally, the SimOTA dynamic label allocation method is applied to One-Category object detection, and the SimOTA-OC dynamic label strategy is proposed to solve the difficult problem of label allocation caused by irregular flying bird objects. In this paper, the performance of the FBOD-SV is validated using experimental datasets of flying bird objects in traction substation surveillance videos. The experimental results show that the FBOD-SV effectively improves the detection performance of flying bird objects in surveillance video. This project is publicly available https://github.com/Ziwei89/FBOD.
@@ I Introduction
BIRD detection plays an extremely important role in ensuring safety and alleviating conflicts between humans and animals and has attracted more and more researchers’ attention in recent years. At present, the method of detecting bird objects by radar is widely used [1, 2], but the radar equipment has the disadvantages of being large in volume, expensive, and having poor visual effects. With the development of computer vision, deep learning, and other technologies, there are more and more studies on using cameras to detect various objects [3, 4, 5, 6]. We are working on the real-time detection of flying birds using surveillance cameras. However, there are three main challenges for detecting flying birds in surveillance video.
@@
    The characteristics of the single frame image of the flying bird object in the surveillance video are not obvious. In most cases, the flying bird object blends into the background, making it challenging to distinguish from the environment, even after careful examination, as shown on the left in Fig. 1LABEL:sub@obvious_bird_fig.
@@
    Most flying bird objects in the surveillance video are small. As Fig. 1LABEL:sub@special_scale_bird_fig demonstrates, when birds are far from the camera, they occupy fewer pixels in the video frame, categorizing them as small objects. Conversely, when birds are closer to the camera, they occupy more pixels and can be considered large objects. However, birds are often far away from the camera, resulting in them being predominantly classified as small objects.
@@
    In most cases, the flying birds in the surveillance video are not symmetrical. For example, in some cases, the body and tail of the bird are divided roughly in half within the bounding box, with half of the tail occupying more background pixels. When the flying bird spreads its wings, the part without wings contains more background pixels, as shown in Fig. 1LABEL:sub@irregularity_scale_bird_fig.
@@
Figure 1: Characteristics of flying bird objects in surveillance videos. (a) On the right, there is a small bird with weak features. Left is a screenshot of the bird on five consecutive frames. (b) Simplified diagram of the distribution of flying birds in surveillance video. Birds are evenly distributed in the surveillance area (A, B, and C areas: birds in area A are large objects, birds in area B are general objects, and birds in area C are small objects). However, the space of area C is much larger than that of areas A and B, so birds are mostly small objects. (c) In the right image, the green box is the bounding box, and the red point is the middle of the bounding box. Due to the shape of the bird object itself, it is not regular in the bounding box in most cases.
@@
Some papers have specifically studied the detection methods of flying bird objects. For example, literature [23] proposed a skeleton-based bird object detection method, which describes the motion information of birds through a set of key postures. However, this method uses the moving object detection method based on VIBE [24] when extracting the contour of the flying bird object. Therefore, this method is unsuitable for situations where the flying bird object is temporarily stationary or the background dynamic changes sharply, which is common in surveillance video.
Another example is Glances and Stare Detection(GSD) [25], which simulates the characteristics of human observation of small moving objects. Firstly, the whole image is glanced at, and the places containing objects are carefully observed. Essentially, this method still relies on single-frame image features (see Algorithm 1 in reference [25]), thus requiring that the single-frame image features of the flying bird object are clear enough. However, the task in our paper needs to deal with the flying bird objects with unobvious features in single frame images, and the method based on single frame image features cannot achieve good results because the features of these objects are easy to lose in the process of feature extraction (see the comparative experiment in Section IV-F1).
@@
Aiming at the problem that the features of the single frame image of the flying bird object in the surveillance video are not obvious, our previous work [26] used the Convolutional Long Short-Term Memory network (ConvLSTM) to aggregate the spatio-temporal information of adjacent multiple frames of flying bird objects at the input of the model to enhance the characteristics of flying bird objects, to improve the detection rate of flying bird objects. However, ConvLSTM suffers from slow runtime and inadequate information aggregation.
Although the characteristics of the single frame image of the flying bird object in the surveillance video are not obvious in some cases, the object can still be found by observing the continuous multiple frames of images, as shown on the right in Fig. 1LABEL:sub@obvious_bird_fig. Therefore, in some cases, the information from other image frames can be used to observe the object on one frame because the bird object has a correlated relationship between frames. In this paper, we take advantage of this correlation relationship and propose a new method to aggregate the information of consecutive multiple frames of images, the Correlation Attention Feature Aggregation (Co-Attention-FA) method, which uses the correlation relationship of flying bird objects on consecutive multiple frames of images to ignore redundant background information and focus on the region containing the flying bird object.
Aiming at the special multi-scale property of flying birds in surveillance videos (most of them belong to small objects), a Flying Bird Object Detection Network (FBOD-Net) is designed, which performs down-sampling followed by up-sampling. A large feature layer that combines fine spatial information and large receptive field information is used to detect flying birds in surveillance videos. The large feature layer has certain advantages for detecting small-scale size bird objects because it has fine spatial information (shallow, high-resolution feature layer) and highly abstract semantic information (deep feature layer). Due to the fusion of large receptive field information, the large feature layer also has the ability to detect large-scale size flying bird objects  [^1] FBOD-Net is trained on datasets dominated by small objects with occasional large objects and can detect large objects. However, the training process may not converge if datasets contain mainly large objects..
Aiming at the irregular characteristics of the bird object in the bounding box in the surveillance video, the SimOTA dynamic label strategy for One-Category object detection (SimOTA-OC) is proposed. When the SimOTA [16] dynamic label allocation method is applied to the bird object detection (one-category object detection), only the IOU between the Ground Truth (GT) box and the predicted box is used to realize the dynamic allocation of labels, which solves the problem of inaccurate label allocation caused by the irregular bird object.
@@
The remainder of this paper is structured as follows: Section II presents work related to this paper. Section III describes the proposed method of detecting flying bird objects in surveillance video in detail. In Section IV, the ablation and comparison experiments of the proposed algorithm (method) are carried out. Section V concludes our work.
@@ II-A Related Work on Feature Aggregation
The related research work of feature aggregation mainly deals with video object detection. When the features of the object to be detected on some video frames are not obvious (appearance change, occlusion, motion blur), the feature information of other adjacent frames can be aggregated to enhance the features of the object on these frames in the feature extraction stage. For example, Zhu Xizhou et al. [27, 28, 29, 30] used optical flow to propagate features, which were then aggregated with the features of the current frame to enhance the features of the object. Haiping Wu et al. [31] proposed a SEquent-Level-Semantic-Aggregation (SELSA) method to enhance the features of the object in the candidate box. Tao Gong et al. [32] aggregated the ROI features of the current frame and the most similar ROI features of other frames to obtain the Temporal ROI features of the object.
In the previous work [26], we pointed out that the above method of feature aggregation, whose aggregation operation is carried out in the intermediate feature layer, is unsuitable for the flying bird object in the surveillance video. Because the appearance features of most of the bird objects in the surveillance video are not particularly rich in any single frame, and the object size is small, the features of the object are easy to lose when extracting the features of a single frame image. After feature extraction, aggregation is easy to introduce some wrong information. Therefore, in the previous work [26], we used the ConvLSTM to aggregate the spatio-temporal information of adjacent multiple frames of flying bird objects at the input of the model to enhance the characteristics of flying bird objects, to improve the detection rate of the flying bird object. However, ConvLSTM suffers from slow runtime and inadequate information aggregation.
@@ III The Proposed FBOD-SV
Fig. 2 shows the summary diagram of the proposed Flying Bird Object Detection method for Surveillance Video (FBOD-SV), which mainly includes four parts, namely, the Co-Attention-FA unit, the FBOD-Net model, the SimOTA-OC label assignment unit, and the model training unit. Among them, the Co-Attention-FA unit aggregates the features of the flying bird object on $n$ consecutive frames of images ($n$=5 as an example in Fig. 2). The FBOD-Net model utilizes aggregated features of the flying bird objects to predict the location information of the flying birds in the middle frame of consecutive $n$ frames (the objects in the middle frame of $n$ consecutive frames have symmetric context information). The SimOTA-OC label assignment unit dynamically allocated positive samples according to the model prediction results during model training. The multi-task loss function is used in the model training unit to train the flying bird object detection model. Next, the Co-Attention-FA unit (III-A), the FBOD-Net (III-B), the SimOTA-OC dynamic label assignment method (III-C), and the loss function (III-D) will be introduced, respectively.
@@ III-A The Co-Attention-FA
The diagram of the Co-Attention-FA module is shown in Fig. 3. Firstly, the consecutive frames of images $\left\{\text{x}_{1},\text{x}_{2},...,\text{x}_{n}\right\}$ are concated into a multi-channel feature matrix X, prepared for the information correlation between frames. Then, the information of consecutive frames is correlated and fused by $3\times 3$ convolution to obtain the feature matrix $\text{F}_{1}$. At this time, $\text{F}_{1}$ already has the correlation information between frames.
@@
Next is a residual structure. Inside the residual structure, $\text{F}_{1}$ is first carried out $3\times 3$ convolution operation to get $\text{F}_{2}$, and the correlation information between frames is used to make $\text{F}_{2}$ focus on the bird object. Then $\text{F}_{2}$ is operated by the Sigmoid function to obtain $\text{F}_{3}$ to make the focusing information explicit. $\text{F}_{4}$ is obtained by dot multiplication of $\text{F}_{3}$ and $\text{F}_{2}$ to realize spatial attention.
On the outside of the residual structure, $\text{F}_{4}$ is superimposed on $\text{F}_{1}$. The purpose of such processing is to highlight the characteristics of the flying bird object while retaining the necessary background information. Finally, the aggregated feature $\text{F}_{6}$ is output by a $3\times 3$ convolution operation.
The flying bird may have different positions and shapes on consecutive video frames. However, the spatial position of the flying bird across $n$ consecutive frames of images is continuous, and its action is coherent [as shown in Fig. 1LABEL:sub@obvious_bird_fig]. Compared to a single frame, the flying bird object exhibits richer features on consecutive frames of images. Therefore, leveraging the characteristics of Co-Attention-FA, we can train the model using backpropagation and gradient descent algorithms to correctly aggregate the features of the flying bird objects across the consecutively $n$ frames of the image.
@@ III-B The FBOD-Net
The flying birds in surveillance videos have special multi-scale properties (i.e., most are small objects), so we adopt a large feature map to predict the flying birds in surveillance videos. However, the shallow large feature maps do not have a large receptive field, which is not conducive to detecting large objects. At the same time, the semantic information of the shallow large feature maps is weak, which is not conducive to object recognition. Therefore, we fuse the information of the deep small feature map of Deep Neural Network (DNN) into the large feature map so that the large feature map has both delicate spatial information and abstract high-level semantic information.
Specifically, we adopt a network structure with down-sampling followed by up-sampling to fully fuse the shallow and deep feature map information, as shown in Fig. 4. Firstly, CSPDarkNet53 [42] was used to extract the down-sampling style features of the feature aggregated image, and a series of feature maps ($\text{C}_{1}$, $\text{C}_{2}$, $\text{C}_{3}$,$\text{C}_{4}$, $\text{C}_{5}$ in Fig. 4) were obtained. Then, the CONVolution (CONV) + Spatial Pyramid Pooling (SPP)+ CONV operation is performed on $\text{C}_{5}$ to obtain the feature map $\text{P}_{5}$. $\text{P}_{5}$ is upsampled (the bilinear interpolation method is used in this paper) and then fused with $\text{C}_{4}$ (a schematic of the fusion operation is shown in Fig. 5) to obtain the feature map $\text{P}_{4}$. The feature layers $\text{P}_{4}$, $\text{P}_{3}$, and $\text{P}_{2}$ perform the same operation as $\text{P}_{5}$ (upsampled and then fused with the shallow feature map of the same scale) and finally obtain the feature map $\text{P}_{1}$. Finally, the feature map $\text{P}_{1}$ undergoes two different convolution operations simultaneously to obtain the feature map of prediction confidence and the feature map of regression flying bird object position.